# Spontaneous time-reversal symmetry breaking by disorder in superconductors

### v1

@@ 2 Introduction
Several superconducting materials have been discovered to manifest evidence of spontaneous time-reversal symmetry breaking (TRSB) that appears only below their critical superconducting transition temperatures [1, 2, 3, 4]. Evidence for such TRSB originates mainly from an enhanced muon-spin relaxation (SR) and/or polar Kerr effect measurements finding a change in the optical polar Kerr angle below . Additional probes able to detect breaking of time-reversal symmetry in the superconducting phase include Josephson interferometry [5], superconducting quantum interference device (SQUID) magnetometry [6], and polarized and small-angle neutron scattering [7]. These experimental methods detect the internal magnetic fields spontaneously generated in the superconducting state. Explanations for the occurrence of such fields from the perspective of clean (homogeneous) physical effects are either in terms of a TRSB non-unitary spin-triplet pairing state or in terms of multi-component superconducting condensates entering a complex TRSB superposition of two superconducting order parameter components. The latter option is particularly natural for superconducting instabilities condensing in two-dimensional (2D) irreducible representations of the associated crystal point group. Complex superpositions of two symmetry-distinct order parameters break TRSB, and generate persistent supercurrent patterns at material edges, dislocations, or around various defect sites, which produce magnetic fields that may not average to zero locally. Such a multi-component scenario for the origin of TRSB below has been extensively discussed for e.g. SrRuO [4, 8, 9, 10, 11, 12], UPt [13, 14], URuSi [15], UTe [16, 17], SrPtAs [18], Re [19], ZrIr [20], LaNiC [21], LaNiGe [22], PrOsSb [23, 24], BaKFeAs [25, 26], FeSeS[27], and more recently also for the kagome superconductors VSb (: K, Rb, Cs) [28, 29, 30] and 4Hb-TaS [31]. For most of the compounds listed here, the evidence for TRSB originates from SR experiments. For further materials displaying TRSB in their superconducting phases, and other types of evidence, we refer to Refs. [1, 2, 3, 4].
@@
At present, the status of the precise superconducting ground state remains controversial for many of the materials listed above. This is the case, for example, for the two materials SrRuO and UTe, where specific heat data features only a single thermodynamic transition that does not split under uniaxial strain, casting considerable doubt on the multi-component nature of their superconducting condensates [32, 33, 34, 35]. Therefore, it appears timely and important to pursue other possibilities for the origin of TRSB detected solely inside the superconducting phase. This motivates the question of what mechanisms there are for TRSB of superconducting condensates composed of a single component order parameter? Barring non-unitary TRSB spin-triplet order, the answer to this question naturally leads to a study of different forms of spatial inhomogeneities and their influence on the local properties of superconductivity. For example, it is known that some superconductors generate magnetic moments below due to an interplay between the gap structure and electronic correlations [36, 37, 38, 39, 40, 41, 40, 42]. As mentioned above, another possibility for TRSB appearing only below includes the generation of localized orbital loops of supercurrents. The latter is well-known to arise near nonmagnetic disorder sites in complex TRSB multi-component condensates [43, 44, 45, 46, 47, 48, 49], but was recently shown to be also present in the strongly disordered regime of single-component superconductors [50, 51]. Likewise, one might expect dislocations and grain boundaries to similarly operate as seeds of localized supercurrents producing sizable internal magnetic fields. Recently, the latter scenario was studied in Refs. [52] and [53] as a possible explanation for reconciling the existence of TRSB and a single specific heat transition in SrRuO. Indeed for this material, edge dislocations are known to be prevalent in many samples [54]. Finally, one might also envision that time-reversal symmetry is already broken above through the existence of small magnetic regions that are either too dilute or rapidly fluctuating to be picked up by most probes. Only through a superconductivity-enhanced coupling, or a sufficient slowing-down of such preexisting magnetism below , can it be detected, e.g. by entering the muon relaxation time window.
@@
Here, motivated by the above-mentioned developments we survey a number of different mechanisms for disorder-induced TRSB in superconductors. We discuss these both in view of recent theoretical and experimental developments of unconventional superconductivity.
@@ 3 Mechanisms of TRSB from disorder in superconductors
In this section we discuss three different mechanisms for disorder-generated TRSB in spin-singlet superconductors. Section 3.1 reviews the generation of magnetic fields from disorder in superconducting phases with TRSB in the bulk or generated near spatial inhomogeneities. Section 3.2 discusses another case where disorder in conjunction with electronic interactions and appropriate superconducting gap structures lead to the generation of static magnetic moments. Due to the required gapping of states for these moments to form, this mechanism is tied to the superconducting phase. Finally, Sec. 3.3 surveys a third case where preexisting magnetic local order (fluctuation moments) in the normal state get enhanced (slowed-down) by superconductivity. Clearly these mechanisms do not exhaust the options for disorder-induced TRSB in superconductors where e.g. cases with substantial spin-orbit coupling may also be of interest [55, 56]. We focus on these three options given recent research along similar lines and for their relevance for some of the superconducting materials mentioned above.
@@ 3.1 Complex combinations of two superconducting condensates
Entering the superconducting state is associated with a breaking of the symmetry of the normal state symmetry group composed of a direct product of crystalline symmetries, gauge symmetry and the time reversal operation [63, 64, 65]. For non-magnetic normal states, the allowed solutions of the gap equation can be characterized by the irreducible representations (irreps) of the associated point group. The leading instability is found from the linearized BCS gap equation as the solution with the largest eigenvalue and thereby also the highest [66, 67, 68, 69]. A state that has broken time-reversal symmetry can arise from a complex combination of superconducting condensates, the breaking of the spin-rotational invariance in non-unitary triplet states, or a breaking of the time-reversal operations itself. Whenever the preferred superconducting instability corresponds to a two-dimensional (2D) irrep, the system may take advantage of the degeneracy between the representations. From a perspective of the minimization of the Ginzburg-Landau free energy, there is the possibility to either form a real combination thereby breaking a crystal symmetry or to form complex superpositions [3, 12, 64]. The latter often maximizes the condensation energy by forming a fully-gapped state and can be shown to be the ground state for a single band setting within a loop expansion [70]. We will not elaborate further on the possibility of real superpositions, but focus instead on complex linear combinations with standard examples as () chiral order () in tetragonal systems and () ordered superconductivity () in hexagonal lattices. Note that by "chiral", we refer to a particular subclass of TRSB superconducting condensates exhibiting phase winding that breaks mirror symmetry. Along the same lines, solutions of one-dimensional (1D) irreps that are accidentally degenerate may also form such complex superpositions. Such cases are fine tuned, but Fermi surfaces with competing nesting tendencies, for example, are prone to near-degeneracy between several symmetry-distinct superconducting instabilities and may therefore more naturally feature such accidental degenerate superpositions. Examples of these include and as discussed in the context of iron-based superconductors and SrRuO (see Fig. 1 B,C), or the so-called state with nontrivial phases on different bands as displayed in the phase diagram in Fig. 1 A, which also arises as a result of frustrated pairing interactions [57, 58, 59]. More recently, accidental degeneracies are also explored for UTe where candidate states include complex spin-triplet superpositions of the form e.g. , or now expressed directly in terms of the 1D irreps, see Fig. 1 D. All the complex superpositions mentioned above break time-reversal symmetry, as evident e.g. from the chosen directionality of the orbital motion of the Cooper pairs.
@@
Even though the above homogeneous superconducting condensates spontaneously break time-reversal symmetry, their unambiguous detection in actual materials is challenging experimentally since the internal magnetic fields produced in these states are suppressed due to the condensate of overlapping Cooper pairs and Meissner screening. However, defects, edges, dislocations, twin boundary interfaces, and other spatial inhomogeneities are known to produce persistent current patterns with associated magnetic fields that do not average to zero [71, 72, 43, 44, 45, 73, 74, 46, 75, 76, 47, 48, 77, 78, 52, 49, 53]. In Fig. 2 we show several examples of disorder- and dislocation-induced current patterns in TRSB superconducting phases. The examples include cases where the homogeneous superconducting phase breaks time-reversal symmetry (Fig. 2) A and B), and a case (Fig. 2) C) where TRSB is only induced locally due to the dislocation, whereas the clean homogeneous superconducting phase preserves time-reversal symmetry, see also Ref. [79]. Naturally, a remaining important question is whether induced fields from this mechanism exhibit the required amplitudes and are voluminous enough to explain experiments of specific materials, or one needs to resort to other explanations for the TRSB signal. We return to this discussion in the final section of the paper.
@@ 3.2.1 Magnetic droplets from impurities
Away from the quantum critical point, the droplet picture has been extensively confirmed for dilute, intentionally substituted defects, primarily by NMR and SR experiments in quasi-1D inorganic chain and ladder compounds [88], as well as in some quasi-2D cuprate materials where impurity substitution into quite clean hosts can be adequately controlled [80]. On the other hand, there is no reason why similar physics should not take place when defects occur naturally in the crystal or film growth process, which is characteristic of many complex oxides, particularly chemically doped ones. The question then naturally arises as to what happens when the inter-defect spacing becomes comparable to the AF correlation length . In calculations for a 2D Hubbard model with disorder and -wave pairing correlations, the overlap of the magnetic droplets in the correlated state leads to interference processes which tend to align the droplets in a state which displays quasi-long range AF order [40, 42, 92, 93, 94]. It seems likely that the results of many experiments which are attributed to "stripe order" in underdoped cuprates may in fact be explained by the quasi-long range order arising from interacting droplets [81, 95]. The creation of 1D structures to lower the kinetic energy via 1D rivers of mobile charge while maintaining some correlation energy is clearly an attractive and simple picture. But creation of such structures also raises the energy of the state around a single impurity, which would prefer to be more globular in shape. One can, within Hubbard-type models on square lattices, answer the question about which mechanism enables the system to lower its energy more efficiently in which situation. This clearly depends on a variety of factors, including doping, correlations, and disorder. By performing simulations, one can - without prejudicing the system - determine which part of the generalized phase diagram corresponds to 1D stripe-like structures, and which to other ground states [96, 97, 81].
Experimentally, manifestations of this physics have been observed in a wide variety of contexts. In Fig. 3 A, we show the well-established oscillations in the Y NMR line in a 1D inorganic chain due to the modulation of the magnetic response locally by a Zn impurity. Figure 3 B shows data from a classic elastic neutron study [89] with frozen magnetic order in a disordered cuprate in zero field, enhanced further by application of a field. The order is apparently quasi-long-ranged since it appears in the neutron scattering intensity near wave vector , yet it occurs in a part of the cuprate phase diagram where true long-range order has disappeared. Finally, in Fig. 3 C, STM reveals in an Fe-based superconductor a spatially modulated density of states pattern due to strain near a twin boundary with a wave vector close the magnetic ordering vector, indicating local impurity-induced magnetism.
@@ 3.2.2 Quasi-long range magnetic order from disorder
Quasi long-range antiferromagnetic order has been reported in many cuprate samples whose or nominal doping puts them outside the long-range AF phase, particularly in the superconducting state at low temperatures, e.g. in LSCO [89, 90, 98, 99, 100, 101, 102, 103]. Characteristic dynamical features of spin glasses have occasionally been observed, but even if these are not always found, the term “spin glass" has become commonplace in cuprates to describe the existence of a phase of short-range frozen magnetism. The size of the spin-glass phase varies significantly with material [103]; whereas in LSCO it extends in zero field out to 15% doping, it is confined to very low doping in the much cleaner YBCO system [104, 105], and BSCCO is in between [103]. Moreover, disorder due to Zn substitution on the Cu sites enhances the size of the spin-glass phase [90]. These empirical facts suggest strongly the picture of disorder-induced short-range magnetism described above. Quasi-long range order occurs when these droplets begin to overlap [40], which can occur without frustration on a square lattice, as shown in Fig. 4 A. This lack of frustration has important implications not only for the quasi-long range () magnetic order, but may also carry a macroscopic net magnetic moment. Since each droplet of staggered order is dominated by the NN spins, it carries a net spin 1/2 [37], and the droplets overlap coherently. While the net moments associated with each droplet do not align ferromagnetically, a fluctuating magnetic state with both short range and longer range (inter-impurity distance) is created (Fig. 4 B). The final magnetization profile appears as a clear peak in the structure factor, as seen in Fig. 4 C, compared to the Lake et al. experiment on underdoped LSCO where intensity at the incommensurate positions is found [89], see in Fig. 4 D. It is this disorder-induced random magnetic landscape that is relevant to our current study as a possible mechanism of TRSB in superconductors.
@@
Similar phenomena have been observed in many other correlated superconductors. Most recently, magnetic clustering was reported by SR in samples of UTe grown by chemical vapor transport with ’s of order 1.6K [107], whereas no magnetic signal was reported in cleaner molten-salt flux grown crystals with ’s of order 2.1K [108].
To summarize this section, simulations of a -wave superconductor in the presence of local Coulomb interactions and nonmagnetic disorder [40, 93, 81, 95] are compatible with the experiments on relatively disordered cuprates like LSCO and BSCCO where a spin-glass phase is observed at low doping in the SC state [103]. These studies explain as well why such a phase is not seen in the cleanest YBCO samples, why in LSCO and BSCCO the spin-glass phase is more robust at underdoping, and why it is enhanced by addition of further disorder. The order-from-disorder many-impurity magnetic state is favored to appear only below when the gap opens. Similar phenomena have been observed in other correlated superconductors.

### v2

@@ 2 Introduction
Several superconducting materials have been discovered to manifest evidence of spontaneous time-reversal symmetry breaking (TRSB) that appears only below their critical superconducting transition temperatures [1, 2, 3, 4]. Evidence for such TRSB originates mainly from an enhanced muon-spin relaxation (SR) and/or polar Kerr effect measurements finding a change in the optical polar Kerr angle below . Additional probes able to detect breaking of time-reversal symmetry in the superconducting phase include Josephson interferometry [5], superconducting quantum interference device (SQUID) magnetometry [6], and polarized and small-angle neutron scattering [7]. These experimental methods detect the internal magnetic fields spontaneously generated in the superconducting state. Explanations for the occurrence of such fields from the perspective of clean (homogeneous) physical effects are either in terms of a TRSB non-unitary spin-triplet pairing state or in terms of multi-component superconducting condensates entering a complex TRSB superposition of two superconducting order parameter components. The latter option is particularly natural for superconducting instabilities condensing in two-dimensional (2D) irreducible representations of the associated crystal point group. Complex superpositions of two symmetry-distinct order parameters break TRSB, and generate persistent supercurrent patterns at material edges, dislocations, or around various defect sites, which produce magnetic fields that may not average to zero locally. Such a multi-component scenario for the origin of TRSB below has been extensively discussed for e.g. Sr2RuO4 [4, 8, 9, 10, 11, 12], UPt3 [13, 14], URu2Si2 [15], UTe2 [16, 17], SrPtAs [18], Re [19], Zr3Ir [20], LaNiC2 [21], LaNiGa2 [22], PrOs4Sb12 [23, 24], Ba1-xKxFe2As2 [25, 26], FeSe1-xSx[27], and more recently also for the kagome superconductors V3Sb5 (: K, Rb, Cs) [28, 29, 30], 4Hb-TaS2 [31], CaPd2Ge2 [32], skutterudites [33, 34] and Na2Cr3As3 [35]. For most of the compounds listed here, the evidence for TRSB originates from SR experiments. For further materials displaying TRSB in their superconducting phases, and other types of evidence, we refer to Refs. [1, 2, 3, 4].
@@
At present, the status of the precise superconducting ground state remains controversial for many of the materials listed above. This is the case, for example, for the two materials Sr2RuO4 and UTe2, where specific heat data features only a single thermodynamic transition that does not split under uniaxial strain, casting considerable doubt on the multi-component nature of their superconducting condensates [36, 37, 38, 39]. Therefore, it appears timely and important to pursue other possibilities for the origin of TRSB detected solely inside the superconducting phase. This motivates the question of what mechanisms there are for TRSB of superconducting condensates composed of a single component order parameter? Barring non-unitary TRSB spin-triplet order, the answer to this question naturally leads to a study of different forms of spatial inhomogeneities and their influence on the local properties of superconductivity. For example, it is known that some superconductors generate magnetic moments below due to an interplay between the gap structure and electronic correlations [40, 41, 42, 43, 44, 45, 44, 46]. As mentioned above, another possibility for TRSB appearing only below includes the generation of localized orbital loops of supercurrents near impurities. The latter is well-known to arise near nonmagnetic disorder sites in complex TRSB multi-component condensates [47, 48, 49, 50, 51, 52, 53, 54], but was recently shown to be also present in the strongly disordered regime of single-component superconductors [55, 56]. Likewise, one might expect dislocations and grain boundaries to similarly operate as seeds of localized supercurrents producing sizable internal magnetic fields. Recently, the latter scenario was studied in Refs. [57] and [58] as a possible explanation for reconciling the existence of TRSB and a single specific heat transition in Sr2RuO4. Indeed for this material, edge dislocations are known to be prevalent in many samples [59]. Finally, one might also envision that time-reversal symmetry is already broken above through the existence of small magnetic regions that are either too dilute or rapidly fluctuating to be picked up by most probes. Only through a superconductivity-enhanced coupling, or a sufficient slowing-down of such preexisting magnetism below , can it be detected, e.g. by entering the muon relaxation time window.
@@
Here, motivated by the above-mentioned developments we survey a number of different mechanisms for inhomogeneity-induced TRSB in superconductors. Our discussion of disorder-induced TRSB refers both to cases where TRSB exists only locally near inhomogeneity sites and where TRS is broken globally, but manifested via induced currents near disorder sites. We discuss these both in view of recent theoretical and experimental developments of unconventional superconductivity.
@@ 3 Mechanisms of TRSB from disorder in superconductors
In this section we discuss three different mechanisms for disorder-generated TRSB in spin-singlet superconductors. Section 3.1 reviews the generation of magnetic fields from disorder in superconducting phases with TRSB in the bulk or generated near spatial inhomogeneities. Section 3.2 discusses another case where disorder in conjunction with electronic interactions and appropriate superconducting gap structures lead to the generation of static magnetic moments. Due to the required gapping of states for these moments to form, this mechanism is tied to the superconducting phase. Finally, Sec. 3.3 surveys a third case where preexisting magnetic local order (fluctuation moments) in the normal state get enhanced (slowed-down) by superconductivity. Clearly these mechanisms do not exhaust the options for disorder-induced TRSB in superconductors where e.g. cases with substantial spin-orbit coupling may also be of interest [60, 61]. We focus on these three options given recent research along similar lines and for their relevance for some of the superconducting materials mentioned above. For a discussion of the possibility of TRSB in homogeneous superconductors originating from conventional electron-phonon coupling, we refer to the review [3] and the references therein. A theoretical proposal for a mechanism has been proposed in Ref. [62] and recent experimental evidence for TRSB in materials where superconductivity is believed to originate from electron-phonon coupling is reported in Refs. [63, 32, 33, 34].
@@ 3.1 Complex combinations of two superconducting condensates
Entering the superconducting state is associated with a breaking of the symmetry of the normal state symmetry group composed of a direct product of crystalline symmetries, gauge symmetry and the time reversal operation [70, 71, 72]. For non-magnetic normal states, the allowed solutions of the gap equation can be characterized by the irreducible representations (irreps) of the associated point group. The leading instability is found from the linearized BCS gap equation as the solution with the largest eigenvalue and thereby also the highest [73, 74, 75, 76]. Broken time-reversal symmetry can arise from a complex combination of superconducting condensates or the breaking of the spin-rotational invariance in non-unitary triplet states. Whenever the preferred superconducting instability corresponds to a two-dimensional (2D) irrep, the system may take advantage of the degeneracy between the representations. From a perspective of the minimization of the Ginzburg-Landau free energy, there is the possibility to either form a real combination thereby breaking a crystal symmetry or to form complex superpositions [3, 12, 71]. The latter often maximizes the condensation energy by forming a fully-gapped state and can be shown to be the ground state for a single band setting within a loop expansion [77]. We will not elaborate further on the possibility of real superpositions, but focus instead on complex linear combinations with standard examples as () chiral order () in tetragonal systems and () ordered superconductivity () in hexagonal lattices. Note that by "chiral", we refer to a particular subclass of TRSB superconducting condensates exhibiting phase winding that breaks mirror symmetries. Along the same lines, solutions of one-dimensional (1D) irreps that are accidentally degenerate may also form such complex superpositions. Such cases are fine tuned, but Fermi surfaces with competing nesting tendencies, for example, are prone to near-degeneracy between several symmetry-distinct superconducting instabilities and may therefore more naturally feature such accidental degenerate superpositions. Examples of these include and as discussed in the context of iron-based superconductors and Sr2RuO4 (see Fig. 1 B,C), or the so-called state with nontrivial phases on different bands as displayed in the phase diagram in Fig. 1 A, which also arises as a result of frustrated pairing interactions which are discussed in view of Fe-based superconductors [64, 65, 66] and other multiband materials[78]. More recently, accidental degeneracies are also explored for UTe2 where candidate states include complex spin-triplet superpositions of the form e.g. , or now expressed directly in terms of the 1D irreps, see Fig. 1 D. All the complex superpositions mentioned above break time-reversal symmetry, as evident e.g. from the chosen directionality of the orbital motion of the Cooper pairs.
@@
Even though the above homogeneous superconducting condensates spontaneously break time-reversal symmetry, their unambiguous detection in actual materials is challenging experimentally since the internal magnetic fields produced in these states are suppressed due to the condensate of overlapping Cooper pairs and Meissner screening. However, defects, edges, dislocations, twin boundary interfaces, and other spatial inhomogeneities are known to produce persistent current patterns with associated magnetic fields that do not average to zero [79, 80, 47, 48, 49, 81, 82, 50, 83, 84, 51, 52, 85, 86, 57, 53, 58]. In Fig. 2 we show several examples of disorder- and dislocation-induced current patterns in TRSB superconducting phases. The examples include cases where the homogeneous superconducting phase breaks time-reversal symmetry (Fig. 2 A and B), and a case (Fig. 2 C) where TRSB is only induced locally due to the dislocation, whereas the clean homogeneous superconducting phase preserves time-reversal symmetry, see also Ref. [87]. Naturally, a remaining important question is whether induced fields from this mechanism exhibit the required amplitudes and are voluminous enough to explain experiments of specific materials, or one needs to resort to other explanations for the TRSB signal. We return to this discussion in the final section of the paper.
@@ 3.2.1 Magnetic droplets from impurities
Away from the quantum critical point, the droplet picture has been extensively confirmed for dilute, intentionally substituted defects, primarily by NMR and SR experiments in quasi-1D inorganic chain and ladder compounds (see Fig. 3 A) [96], as well as in some quasi-2D cuprate materials where impurity substitution into quite clean hosts can be adequately controlled [88]. On the other hand, there is no reason why similar physics should not take place when defects occur naturally in the crystal or film growth process, which is characteristic of many complex oxides, particularly chemically doped ones. The question then naturally arises as to what happens when the inter-defect spacing becomes comparable to the AF correlation length . In calculations for a 2D Hubbard model with disorder and -wave pairing correlations, the overlap of the magnetic droplets in the correlated state leads to interference processes which tend to align the droplets in a state which displays quasi-long range AF order [44, 46, 100, 101, 102]. It seems likely that the results of many experiments which are attributed to "stripe order" in underdoped cuprates may in fact be explained by the quasi-long range order arising from interacting droplets [89, 103]. The creation of 1D structures to lower the kinetic energy via 1D rivers of mobile charge while maintaining some correlation energy is clearly an attractive and simple picture. But creation of such structures also raises the energy of the state around a single impurity, which would prefer to be more globular in shape. One can, within Hubbard-type models on square lattices, answer the question about which mechanism enables the system to lower its energy more efficiently in which situation. This clearly depends on a variety of factors, including doping, correlations, and disorder. By performing simulations, one can - without prejudicing the system - determine which part of the generalized phase diagram corresponds to 1D stripe-like structures, and which to other ground states [104, 105, 89].
Experimentally, manifestations of this physics have been observed in a wide variety of contexts. In Fig. 3 A, we show the well-established oscillations in the Y NMR line in a 1D inorganic chain due to the modulation of the magnetic response locally by a Zn impurity. Figure 3 B shows data from an elastic neutron study [97] (see also [106, 107, 108]) with frozen magnetic order in a disordered cuprate in zero field, enhanced further by application of a field. The order is apparently quasi-long-ranged since it appears in the neutron scattering intensity near wave vector , yet it occurs in a part of the cuprate phase diagram where true long-range order has disappeared. Finally, in Fig. 3 C, STM reveals in an Fe-based superconductor a spatially modulated density of states pattern due to strain near a twin boundary with a wave vector close the magnetic ordering vector, indicating local impurity-induced magnetism.
@@ 3.2.2 Quasi-long range magnetic order from disorder
Quasi long-range antiferromagnetic order has been reported in many cuprate samples whose or nominal doping puts them outside the long-range AF phase, particularly in the superconducting state at low temperatures, e.g. in LSCO [106, 98, 109, 97, 107, 110, 111, 112]. Characteristic dynamical features of spin glasses have occasionally been observed, but even if these are not always found, the term “spin glass" has become commonplace in cuprates to describe the existence of a phase of short-range frozen magnetism. The size of the spin-glass phase varies significantly with material [112]; whereas in LSCO it extends in zero field out to 15% doping, it is confined to very low doping in the much cleaner YBCO system [113, 114], and BSCCO is in-between [112]. Moreover, disorder due to Zn substitution on the Cu sites enhances the size of the spin-glass phase [98]. These empirical facts suggest strongly the picture of disorder-induced short-range magnetism described above. Quasi-long range order occurs when these droplets begin to overlap [44], which can occur without frustration on a square lattice, as shown in Fig. 4 A. This lack of frustration has important implications not only for the quasi-long range () magnetic order, but may also carry a macroscopic net magnetic moment. Since each droplet of staggered order is dominated by the NN spins, it carries a net spin 1/2 [41], and the droplets overlap coherently. While the net moments associated with each droplet do not align ferromagnetically, a fluctuating magnetic state with both short range and longer range (inter-impurity distance) is created (Fig. 4 B). The final magnetization profile appears as a clear peak in the structure factor, as seen in Fig. 4 C, compared to the Lake et al. experiment on underdoped LSCO where intensity at the incommensurate positions is found [106], as shown schematically in Fig. 4 D. It is this disorder-induced random magnetic landscape that is relevant to our current study as a possible mechanism of TRSB in superconductors.
@@
Similar phenomena have been observed in many other correlated superconductors. Most recently, magnetic clustering was reported by SR in samples of UTe2 grown by chemical vapor transport with ’s of order 1.6K [116], whereas no magnetic signal was reported in cleaner molten-salt flux grown crystals with ’s of order 2.1K [117].
To summarize this section, simulations of a -wave superconductor in the presence of local Coulomb interactions and nonmagnetic disorder [44, 101, 89, 103] are compatible with the experiments on relatively disordered cuprates like LSCO and BSCCO where a spin-glass phase is observed at low doping in the SC state [112]. These studies explain as well why such a phase is not seen in the cleanest YBCO samples, why in LSCO and BSCCO the spin-glass phase is more robust at underdoping, and why it is enhanced by addition of further disorder. The order-from-disorder many-impurity magnetic state is favored to appear only below when the gap opens. Similar phenomena have been observed in other correlated superconductors [118, 119, 120, 91].